# A real-time digital twin of azimuthal thermoacoustic instabilities

### v1

@@ Abstract
When they occur, azimuthal thermoacoustic oscillations can detrimentally affect the safe operation of gas turbines and aeroengines. We develop a real-time digital twin of azimuthal thermoacoustics of a hydrogen-based annular combustor. The digital twin seamlessly combines two sources of information about the system (i) a physics-based low-order model; and (ii) raw and sparse experimental data from microphones, which contain both aleatoric noise and turbulent fluctuations. First, we derive a low-order thermoacoustic model for azimuthal instabilities, which is deterministic. Second, we propose a real-time data assimilation framework to infer the acoustic pressure, the physical parameters, and the model and measurement biases simultaneously. This is the bias-regularized ensemble Kalman filter (r-EnKF), for which we find an analytical solution that solves the optimization problem. Third, we propose a reservoir computer, which infers both the model bias and measurement bias to close the assimilation equations. Fourth, we propose a real-time digital twin of the azimuthal thermoacoustic dynamics of a laboratory hydrogen-based annular combustor for a variety of equivalence ratios. We find that the real-time digital twin (i) autonomously predicts azimuthal dynamics, in contrast to bias-unregularized methods; (ii) uncovers the physical acoustic pressure from the raw data, i.e., it acts as a physics-based filter; (iii) is a time-varying parameter system, which generalizes existing models that have constant parameters, and capture only slow-varying variables. The digital twin generalizes to all equivalence ratios, which bridges the gap of existing models. This work opens new opportunities for real-time digital twinning of multi-physics problems.
@@ 1 Introduction
The intricate linear and nonlinear dynamics of azimuthal thermoacoustic oscillations spurred interest in physics-based low-order modelling and control (e.g., Morgans and Stow, 2007; Illingworth and Morgans, 2010; Humbert et al., 2023b). The azimuthal dynamics can be qualitatively described by a one-dimensional wave-like equation (Noiray et al., 2011; Ghirardo and Juniper, 2013; Bauerheim et al., 2014; Bothien et al., 2015; Yang et al., 2019), which can also include a stochastic forcing term to model the effect of the turbulent fluctuations and noise (Noiray and Schuermans, 2013a; Orchini et al., 2020). In the frequency domain, azimuthal oscillations were investigated with eigenvalue sensitivity Magri et al. (2016a, b); Mensah et al. (2019); Orchini et al. (2020), who found that traditional eigenvalue sensitivity analysis needed to be extended to tackle degenerate pairs of azimuthal modes, as reviewed in Magri et al. (2023). A review of azimuthal thermoacoustic modelling can be found in Bauerheim et al. (2014). In the time domain, the typical approach is to develop models that describe the acoustic state, which is also referred to as the slow-varying variable approach based on quaternions (e.g., Ghirardo and Bothien, 2018) or as the generalized Bloch sphere representation (Magri et al., 2023). In this approach, all fast-varying dynamics, such as aleatoric noise and turbulent fluctuations, are filtered out of the data and modelled as a stochastic forcing term. The equations model the amplitude of the acoustic pressure envelope, the temporal phase drift, and the angles defining rotational and reflectional symmetry breaking (Ghirardo and Bothien, 2018). Faure-Beaulieu and Noiray (2020) introduced the slow-varying variables into an stochastic wave equation, which was averaged in space and time to obtain coupled Langevin equations. The low-order model parameters were estimated via Langevin regression, which is a regression method used in turbulent environments (e.g., Noiray, 2017; Boujo et al., 2020; Callaham et al., 2021). Slow-varying variable models provide qualitatively accurate representations of azimuthal oscillations. However, they are not suitable for real-time applications, in which we need models that can receive data from sensors as raw inputs. Therefore, this work focus on models describing the fast-varying quantities, which model the evolution of the modal amplitudes through coupled Van der Pol oscillators (e.g., Noiray et al., 2011). Nonetheless, the discussed low-order modelling methods are offline, i.e., they infer the model parameters from the data in a post-processing stage and they identify one parameter at a time (i.e., they do not estimate all the parameters in one computation). This means that the physical parameters of the literature are not necessarily optimal.
@@ 1.1 Objectives and structure
The objective of this work is to develop a real-time digital twin of azimuthal thermoacoustic instabilities. Real-time digital twins are adaptive models, which are designed to predict the behaviour of their physical counterpart by assimilating data from sensors when it becomes available. We need four ingredients to design a real-time digital twin (i) data from sensors (from noisy and sparse microphones); (ii) a qualitative low-order model, which should capture the fast-varying acoustic dynamics rather than the slow-varying states; (iii) an estimator of both the model bias and measurement bias, (iv) a statistical data assimilation method that optimally combines the data and the model on the fly (real time) to infer the physical states and parameters. §2 describes the available experimental data, as well as the physics-based low-order model used to describe thermoacoustic problem. The real-time assimilation framework is detailed in §3. The proposed ESN is described in §4. §5 shows the results of the real-time digital twin, and compares the performance of the bias-unregularized EnKF with the bias-regularized r-EnKF. §6 ends the paper.
@@ 2.1 Experimental setup and data
Figure 2: Pressure at the four measurement locations with equivalence ratio $\Phi=0.5125$ (thermoacoustically unstable configuration). (a) Raw experimental data, i.e., the observables, and (b) post-processed data, which is briefly referred to as the truth. (I) Time series of the fast-varying oscillations, and (II) histogram of the acoustic pressure in the time window for data assimilation, $t\in[0.5,0.85]$. The horizontal lines in (II) indicate the mean of the histograms. The thermoacoustic limit cycle is a standing acoustic mode with a nodal line near $\theta=120^{\circ}$ (mono-modal histogram).
Figure 2 shows the time series and histograms of the raw and post-processed experimental data at $\Phi=0.5125$ (Figure 2a,b). The pressure measurements have a mean value that is different from zero because of large-scale flow structures, which are not correlated to the thermoacoustic dynamics. The non-zero mean pressure is referred to as “measurement bias”, as further explained in §3.1. In a digital twin that is coupled with sensors’ measurements in real time, the data are the raw pressure signals that are assimilated into the low-order model (Figure 2a). Because the raw pressure contains aleatoric noise, measurement bias, and turbulent flow fluctuations, we refer to the post-processed data (Figure 2b) as the “presumed ground truth” or “presumed acoustic state”, that is, the physical acoustic state that is the presumed target prediction.
@@ 3.1 Aleatoric and epistemic uncertainties
${\bm{\mathit{\phi}}}+{\bm{\mathit{\epsilon}}}_{\phi}={\bm{\mathit{\phi}}}^{\dagger},\quad{\bm{\mathit{\alpha}}}+{\bm{\mathit{\epsilon}}}_{\alpha}={\bm{\mathit{\alpha}}}^{\dagger},$ (8)
where $\dagger$ indicates the true quantity (which is unknown). The aleatoric uncertainties are modelled as Gaussian distributions
@@
where $\mathcal{N}({\bm{\mathit{0}}},\bm{\mathsf{C}})$ is a normal distribution with zero mean and covariance $\bm{\mathsf{C}}$. Second, we discuss model biases, which are epistemic uncertainties (Nóvoa et al., 2024). The model bias may arise from modelling assumptions and approximations in the operators $\mathcal{F}$ and $\mathcal{M}$. It is unknown a priori (colloquially, an “unknown unknown” (Nóvoa et al., 2024)). The bias is
${\bm{\mathit{b}}}={\bm{\mathit{d}}}^{\dagger}-\mathbb{E}({\bm{\mathit{q}}}),$ (10)
which is the expected difference between the true observable and the model observable. (If the model is unbiased, ${\bm{\mathit{d}}}^{\dagger}=\mathbb{E}({\bm{\mathit{q}}})$.) Hence, the bias-corrected model observable is
${\bm{\mathit{y}}}={\bm{\mathit{q}}}+{\bm{\mathit{b}}}+{\bm{\mathit{\epsilon}}}_{q},$ (11)
where ${\bm{\mathit{\epsilon}}}_{q}\sim\mathcal{N}({\bm{\mathit{0}}},\bm{\mathsf{C}}_{qq})$. The final model equations, which define the first source of information on the system, are
@@
The set of equations (14) is not closed because we need a model for the model bias (§4).
The second source of information on the system is the data measured by the sensors. Experimental data are affected by both aleatoric noise and measurement biases (§2.1). The measurements at a time instant are
${\bm{\mathit{d}}}+{\bm{\mathit{b}}}_{d}+{\bm{\mathit{\epsilon}}}_{d}={\bm{\mathit{d}}}^{\dagger},$ (15)
where the measurement bias ${\bm{\mathit{b}}}_{d}$ is defined as
${\bm{\mathit{b}}}_{d}=\mathbb{E}({\bm{\mathit{d}}})-{\bm{\mathit{d}}}^{\dagger}.$ (16)
In the problem under investigation, ${\bm{\mathit{d}}}$ is the raw acoustic data, ${\bm{\mathit{d}}}^{\dagger}$ is is the post-processed data, and ${\bm{\mathit{b}}}_{d}$ is the non-zero mean of the raw data (see §2.1). The aleatoric noise affects the measurement as ${\bm{\mathit{\epsilon}}}_{d}\sim\mathcal{N}({\bm{\mathit{0}}},\bm{\mathsf{C}}_{dd})$. For simplicity, we assume that the measurement errors are statistically independent, i.e., $\bm{\mathsf{C}}_{dd}$ is a diagonal matrix with diagonal ${\bm{\mathit{\epsilon}}}_{d}$.
Both model and measurement biases are unknown a priori. To infer them, we analyse the residuals between the forecast and the observations, which are also known as innovations (Dee and Todling, 2000; Haimberger, 2007)
${\bm{\mathit{i}}}={\bm{\mathit{d}}}-{\bm{\mathit{q}}}.$ (17)
Hence, the expected value of the innovation is
$\mathbb{E}({\bm{\mathit{i}}})={\bm{\mathit{b}}}_{d}+{\bm{\mathit{b}}},$ (18)
i.e., the expected innovation is the sum of the measurement and model biases. The relation between the innovations ${\bm{\mathit{i}}}$ and the biases ${\bm{\mathit{b}}}$ and ${\bm{\mathit{b}}}_{d}$ will be essential for the design of the bias estimator in §4.
@@ 3.2 Augmented state-space formulation
$\left\{\begin{array}[]{rcl}\mathrm{d}\,\begin{bmatrix}{\bm{\mathit{\phi}}}\\ {\bm{\mathit{\alpha}}}\\ {\bm{\mathit{q}}}\end{bmatrix}&=&\begin{bmatrix}\mathcal{F}({\bm{\mathit{\phi}}}+{\bm{\mathit{\epsilon}}}_{\phi},{\bm{\mathit{\alpha}}}+{\bm{\mathit{\epsilon}}}_{\alpha})\\ {\bm{\mathit{0}}}_{N_{\alpha}}\\ {\bm{\mathit{0}}}_{N_{q}}\end{bmatrix}{\mathrm{d}\,t}\\[20.00003pt] {\bm{\mathit{y}}}&=&{\bm{\mathit{q}}}+{\bm{\mathit{b}}}+{\bm{\mathit{\epsilon}}}_{q}\end{array}\right.\quad\leftrightarrow\quad\left\{\begin{array}[]{rcl}\mathrm{d}\,{\bm{\mathit{\psi}}}&=&\bm{\mathsf{F}}\left({\bm{\mathit{\psi}}}+{\bm{\mathit{\epsilon}}}_{\psi}\right){\mathrm{d}\,t}\\[10.00002pt] {\bm{\mathit{y}}}&=&\bm{\mathsf{M}}{\bm{\mathit{\psi}}}+{\bm{\mathit{b}}}+{\bm{\mathit{\epsilon}}}_{q}\end{array}\right.$ (23)
where $\bm{\mathsf{F}}({\bm{\mathit{\psi}}})$ and ${\bm{\mathit{\epsilon}}}_{\psi}$ are the augmented nonlinear operator and aleatoric uncertainties, respectively; $\bm{\mathsf{M}}=\left[\bm{\mathsf{0}}~{}\big{|}~{}\mathbb{I}_{N_{q}}\right]$ is the linear measurement operator, which consists of the vertical concatenation of a matrix of zeros, $\bm{\mathsf{0}}\in\mathbb{R}^{N_{q}\times(N_{\phi}+N_{\alpha})}$, and the identity matrix, $\mathbb{I}_{N_{q}}\in\mathbb{R}^{N_{q}\times N_{q}}$; and ${\bm{\mathit{0}}}_{N_{\alpha}}$ and ${\bm{\mathit{0}}}_{N_{q}}$ are vectors of zeros (because the parameters are constant in time, and ${\bm{\mathit{q}}}$ is not integrated in time but it is only computed at the analysis step).

### v2

@@ Abstract
When they occur, azimuthal thermoacoustic oscillations can detrimentally affect the safe operation of gas turbines and aeroengines. We develop a real-time digital twin of azimuthal thermoacoustics of a hydrogen-based annular combustor. The digital twin seamlessly combines two sources of information about the system (i) a physics-based low-order model; and (ii) raw and sparse experimental data from microphones, which contain both aleatoric noise and turbulent fluctuations. First, we derive a low-order thermoacoustic model for azimuthal instabilities, which is deterministic. Second, we propose a real-time data assimilation framework to infer the acoustic pressure, the physical parameters, and the model bias and measurement shift simultaneously. This is the bias-regularized ensemble Kalman filter (r-EnKF), for which we find an analytical solution that solves the optimization problem. Third, we propose a reservoir computer, which infers both the model bias and measurement shift to close the assimilation equations. Fourth, we propose a real-time digital twin of the azimuthal thermoacoustic dynamics of a laboratory hydrogen-based annular combustor for a variety of equivalence ratios. We find that the real-time digital twin (i) autonomously predicts azimuthal dynamics, in contrast to bias-unregularized methods; (ii) uncovers the physical acoustic pressure from the raw data, i.e., it acts as a physics-based filter; (iii) is a time-varying parameter system, which generalizes existing models that have constant parameters, and capture only slow-varying variables. The digital twin generalizes to all equivalence ratios, which bridges the gap of existing models. This work opens new opportunities for real-time digital twinning of multi-physics problems.
@@ 1 Introduction
The intricate linear and nonlinear dynamics of azimuthal thermoacoustic oscillations spurred interest in physics-based low-order modelling and control (e.g., Morgans and Stow, 2007; Illingworth and Morgans, 2010; Humbert et al., 2023b). The azimuthal dynamics can be qualitatively described by a one-dimensional wave-like equation (Noiray et al., 2011; Ghirardo and Juniper, 2013; Bauerheim et al., 2014; Bothien et al., 2015; Yang et al., 2019), which can also include a stochastic forcing term to model the effect of the turbulent fluctuations and noise (Noiray and Schuermans, 2013a; Orchini et al., 2020). In the frequency domain, azimuthal oscillations were investigated with eigenvalue sensitivity Magri et al. (2016a, b); Mensah et al. (2019); Orchini et al. (2020), who found that traditional eigenvalue sensitivity analysis needed to be extended to tackle degenerate pairs of azimuthal modes, as reviewed in Magri et al. (2023). A review of azimuthal thermoacoustic modelling can be found in Bauerheim et al. (2014). In the time domain, the typical approach is to develop models that describe the acoustic state, which is also referred to as the slow-varying variable approach based on quaternions (e.g., Ghirardo and Bothien, 2018) or as the generalized Bloch sphere representation (Magri et al., 2023). In this approach, all fast-varying dynamics, such as aleatoric noise and turbulent fluctuations, are filtered out of the data and modelled as a stochastic forcing term. The equations model the amplitude of the acoustic pressure envelope, the temporal phase drift, and the angles defining rotational and reflectional symmetry breaking (Ghirardo and Bothien, 2018). Faure-Beaulieu and Noiray (2020) introduced the slow-varying variables into an stochastic wave equation, which was averaged in space and time to obtain coupled Langevin equations. The low-order model parameters were estimated via Langevin regression, which is a regression method used in turbulent environments (e.g., Siegert et al., 1998; Noiray, 2017; Boujo et al., 2020; Callaham et al., 2021). Slow-varying variable models provide qualitatively accurate representations of azimuthal oscillations. However, they are not suitable for real-time applications, in which we need models that can receive data from sensors as raw inputs. Therefore, this work focus on models describing the fast-varying quantities, which model the evolution of the modal amplitudes through coupled Van der Pol oscillators (e.g., Noiray et al., 2011). Nonetheless, the discussed low-order modelling methods are offline, i.e., they infer the model parameters from the data in a post-processing stage and they identify one parameter at a time (i.e., they do not estimate all the parameters in one computation). This means that the physical parameters of the literature are not necessarily optimal.
@@ 1.1 Objectives and structure
The objective of this work is to develop a real-time digital twin of azimuthal thermoacoustic instabilities. Real-time digital twins are adaptive models, which are designed to predict the behaviour of their physical counterpart by assimilating data from sensors when it becomes available. We need four ingredients to design a real-time digital twin (i) data from sensors (from noisy and sparse microphones); (ii) a qualitative low-order model, which should capture the fast-varying acoustic dynamics rather than the slow-varying states; (iii) an estimator of both the model bias and measurement shift, (iv) a statistical data assimilation method that optimally combines the data and the model on the fly (real time) to infer the physical states and parameters. §2 describes the available experimental data, as well as the physics-based low-order model used to describe thermoacoustic problem. The real-time assimilation framework is detailed in §3. The proposed ESN is described in §4. §6 shows the results of the real-time digital twin, and compares the performance of the bias-unregularized EnKF with the bias-regularized r-EnKF. §7 ends the paper.
@@ 2.1 Experimental setup and data
Figure 2: Pressure at the four measurement locations with equivalence ratio $\Phi=0.5125$ (thermoacoustically unstable configuration). (a) Raw experimental data, i.e., the observables, and (b) post-processed data, which is briefly referred to as the presumed truth. (I) Time series of the fast-varying oscillations, (II) histogram of the acoustic pressure in the time window for data assimilation, $t\in[0.5,0.85]$, and (III) comparison of the power spectral density of the raw data (light colour) and the presumed truth (dark colour). The horizontal lines in (II) indicate the mean of the histograms. The thermoacoustic limit cycle is a standing acoustic mode with a nodal line near $\theta=120^{\circ}$ (mono-modal histogram).
Figure 2 shows the time series, histograms and power spectral density (PSD) of the raw and post-processed experimental data at $\Phi=0.5125$. The PSD of the post-processed data is obtained after applying a Butterworth band-pass filter to the raw data around the frequency of instability, which approximately isolates the frequencies between 1050 and 1150 Hz (Figure 2III). The pressure measurements have a mean value that is different from zero because of large-scale flow structures, which are not correlated to the thermoacoustic dynamics. The non-zero mean pressure is referred to as “measurement shift”, as further explained in §3.1. In a digital twin that is coupled with sensors’ measurements in real time, the data are the raw pressure signals that are assimilated into the low-order model (Figure 2a). Because the raw pressure contains aleatoric noise, measurement shift, and turbulent flow fluctuations, we refer to the post-processed data (Figure 2b) as the “presumed ground truth” or “presumed acoustic state”, that is, the physical acoustic state that is the presumed target prediction.
@@ 3 Real-time data assimilation
Figure 3: Schematic of the proposed digital twin framework. (a) The physical and digital systems evolve in time (left to right) independently. The digital system is composed of a physical model with uncertain state and parameters, and an estimator of the model bias and innovations (from which the measurement shift is estimated). When sensor data (crosses) become available, they are combined with the estimates from the digital twin using the regularized ensemble Kalman filter (r-EnKF) to update the digital system. (b) Diagram of the r-EnKF update performed sequentially every time that data become available.
@@ 3.1 Aleatoric and epistemic uncertainties
${\bm{\mathit{\phi}}}+{\bm{\mathit{\epsilon}}}_{\phi}={\bm{\mathit{\phi}}}^{\mathrm{t}},\quad{\bm{\mathit{\alpha}}}+{\bm{\mathit{\epsilon}}}_{\alpha}={\bm{\mathit{\alpha}}}^{\mathrm{t}},$ (8)
where $\mathrm{t}$ indicates the true quantity (which is unknown). The aleatoric uncertainties are modelled as Gaussian distributions
@@
where $\mathcal{N}({\bm{\mathit{0}}},\bm{\mathsf{C}})$ is a normal distribution with zero mean and covariance $\bm{\mathsf{C}}$. Second, we discuss model biases, which are epistemic uncertainties (Nóvoa et al., 2024). The model bias may arise from modelling assumptions and approximations in the operators $\mathcal{F}$ and $\mathcal{M}$. The true bias of a model is defined as
${{\bm{\mathit{b}}}^{\mathrm{t}}={\bm{\mathit{d}}}^{\mathrm{t}}}-\mathbb{E}({\bm{\mathit{q}}}),$ (10)
which is the expected difference between the true observable and the model observable. (If the model is unbiased, ${{\bm{\mathit{d}}}^{\mathrm{t}}}=\mathbb{E}({\bm{\mathit{q}}})$.) The true bias is unknown (colloquially, an “unknown unknown” (Nóvoa et al., 2024)) because we do not have access to the ground truth. Therefore, we need to estimate it from the information available from the data. We referred to as the presumed true bias as ${\bm{\mathit{b}}}^{\dagger}={\bm{\mathit{d}}}^{\mathrm{t}}-\mathbb{E}({\bm{\mathit{q}}})$, and to the bias estimated from a model as ${\bm{\mathit{b}}}$. With this, the model equations, which define the first source of information on the system, are
@@
where ${\bm{\mathit{y}}}$ is the bias-corrected model observable, and ${\bm{\mathit{\epsilon}}}_{q}\sim\mathcal{N}({\bm{\mathit{0}}},\bm{\mathsf{C}}_{qq})$. The set of equations (13) is not closed until we define an estimator for the model bias (§4).
The second source of information on the system is the data measured by the sensors. Experimental data are affected by both aleatoric noise and measurement shifts (§2.1). The true measurement shift is
${\bm{\mathit{b}}}_{d}^{\mathrm{t}}=\mathbb{E}({\bm{\mathit{d}}})-{\bm{\mathit{d}}}^{\mathrm{t}}.$ (14)
Because the ground truth is not known, the best estimate on the model bias is the presumed true measurement shift ${\bm{\mathit{b}}}_{d}^{\dagger}={\bm{\mathit{d}}}^{\dagger}-\mathbb{E}({\bm{\mathit{d}}})$ and the estimated measurement shift is ${\bm{\mathit{b}}}_{d}$. With this, we define the measurements at a time instant as
${\bm{\mathit{d}}}+{\bm{\mathit{b}}}_{d}+{\bm{\mathit{\epsilon}}}_{d}={\bm{\mathit{d}}}^{\mathrm{t}}.$ (15)
In the problem under investigation, ${\bm{\mathit{d}}}$ is the raw acoustic data, ${\bm{\mathit{d}}}^{\dagger}$ is is the post-processed data, and ${\bm{\mathit{b}}}_{d}$ is the non-zero mean of the raw data (see §2.1). The aleatoric noise affects the measurement as ${\bm{\mathit{\epsilon}}}_{d}\sim\mathcal{N}({\bm{\mathit{0}}},\bm{\mathsf{C}}_{dd})$. For simplicity, we assume that the measurement errors are statistically independent, i.e., $\bm{\mathsf{C}}_{dd}$ is a diagonal matrix with identical diagonal entries $\sigma_{d}$.
Both the model bias and measurement shift are unknown a priori. To infer them, we analyse the residuals between the forecast and the observations, which are also known as innovations (Dee and Todling, 2000; Haimberger, 2007), which are defined as
${\bm{\mathit{i}}}={\bm{\mathit{d}}}-{\bm{\mathit{q}}}\quad\Rightarrow\quad\mathbb{E}({\bm{\mathit{i}}})={\bm{\mathit{b}}}_{d}+{\bm{\mathit{b}}}.$ (16)
Therefore, the expected innovation is the sum of the measurement and model biases as defined in (10) and (14). The relation between the innovations ${\bm{\mathit{i}}}$ and the biases ${\bm{\mathit{b}}}$ and ${\bm{\mathit{b}}}_{d}$ will be essential for the design of the bias estimator in §4.
Table 1: Summary of the terminology.
Symbol Description
${\bm{\mathit{d}}}^{\mathrm{t}}$ Ground truth (unknown).
${\bm{\mathit{d}}}^{\dagger}$ Presumed truth, obtained by applying a Butterworth filter to the raw data.
${\bm{\mathit{d}}}$ Raw experimental data from microphones (15).
${\bm{\mathit{b}}}^{\mathrm{t}},{\bm{\mathit{b}}}_{d}^{\mathrm{t}}$ True model bias and measurement shift (unknown). Eqs. (10) and (14).
${\bm{\mathit{b}}}^{\dagger},{\bm{\mathit{b}}}_{d}^{\dagger}$ Presumed true model bias and measurement shift.
${\bm{\mathit{b}}}^{\,},{\bm{\mathit{b}}}_{d}$ Estimates of the model bias and measurement shift (§ 4).
@@ 3.2 Augmented state-space formulation
$\left\{\begin{array}[]{rcl}\mathrm{d}\,\begin{bmatrix}{\bm{\mathit{\phi}}}\\ {\bm{\mathit{\alpha}}}\\ {\bm{\mathit{q}}}\end{bmatrix}&=&\begin{bmatrix}\mathcal{F}({\bm{\mathit{\phi}}}+{\bm{\mathit{\epsilon}}}_{\phi},{\bm{\mathit{\alpha}}}+{\bm{\mathit{\epsilon}}}_{\alpha})\\ {\bm{\mathit{0}}}_{N_{\alpha}}\\ {\bm{\mathit{0}}}_{N_{q}}\end{bmatrix}{\mathrm{d}\,t}\\[20.00003pt] {\bm{\mathit{y}}}&=&{\bm{\mathit{q}}}+{\bm{\mathit{b}}}+{\bm{\mathit{\epsilon}}}_{q}\end{array}\right.\;\leftrightarrow\;\left\{\begin{array}[]{rcl}\mathrm{d}\,{\bm{\mathit{\psi}}}&=&\bm{\mathsf{F}}\left({\bm{\mathit{\psi}}}+{\bm{\mathit{\epsilon}}}_{\psi}\right){\mathrm{d}\,t}\\[10.00002pt] {\bm{\mathit{y}}}&=&\bm{\mathsf{M}}{\bm{\mathit{\psi}}}+{\bm{\mathit{b}}}+{\bm{\mathit{\epsilon}}}_{q}\end{array}\right.{,\;\forall\,t\neq{t_{d}}}$ (21)
where $t_{d}$ are the times when the assimilation is performed, $\bm{\mathsf{F}}({\bm{\mathit{\psi}}})$ and ${\bm{\mathit{\epsilon}}}_{\psi}$ are the augmented nonlinear operator and aleatoric uncertainties, respectively; $\bm{\mathsf{M}}=\left[\bm{\mathsf{0}}~{}\big{|}~{}\mathbb{I}_{N_{q}}\right]$ is the linear measurement operator, which consists of the vertical concatenation of a matrix of zeros, $\bm{\mathsf{0}}\in\mathbb{R}^{N_{q}\times(N_{\phi}+N_{\alpha})}$, and the identity matrix, $\mathbb{I}_{N_{q}}\in\mathbb{R}^{N_{q}\times N_{q}}$; and ${\bm{\mathit{0}}}_{N_{\alpha}}$ and ${\bm{\mathit{0}}}_{N_{q}}$ are vectors of zeros (because the parameters are constant in time, and ${\bm{\mathit{q}}}$ is not integrated in time but it is only computed at the analysis step).